OpenStudy (anonymous):

(2n+7)/(n-1) - (8n)/(n+5)

OpenStudy (anonymous):

LCD is (n-1)(n+5) \[[(n-1)(n-5)](\frac{2n+7}{n-1}-\frac{8n}{n+5})\] \[\frac{(2n+7)(n-5)-8n(n-1)}{(n-1)(n+5)}\] \[\frac{2n^2-3n-35-8n^2+8n}{(n-1)(n+5)}\] \[\frac{-6n^2-5n-35}{(n-1)(n+5)}\] \[\frac{-1(6n^2+5n+35)}{(n-1)(n+5)}\] This can't be factored

OpenStudy (anonymous):

Did you write the question properly?

OpenStudy (anonymous):

Yes. I didn't make any typos. In case you are a little confused, though, I'll post a drawing of what it says.

OpenStudy (anonymous):